# Monitored long-range interacting systems: spin-wave theory for quantum trajectories

### v1

@@ Appendix D Generalization to spin-boson systems
In this appendix, we demonstrate a possible extension of our spin-wave theory, which allows the investigation of spin-boson systems. We adopt the same Gaussian approximation for the bosonic mode as for the bosonized spins, such that the state of the entire system can be characterized by the first and second moments, requiring a quadratic number of variables to specify the Gaussian variational ansatz in the general case. For illustration purposes, let us consider a driven-dissipative Tavis-Cummings model with long-range spin interactions. The spins are collectively driven and resonantly coupled to a single-mode cavity (with annihilation operator $\hat{a}$ for the cavity bosonic mode), and the cavity undergoes single-photon dissipation. The Hamiltonian can be written as follows,
@@
$\hat{H}=$ $\leavevmode\nobreak\ \omega\hat{S}^{x}+\dfrac{2sJ}{\mathcal{N}}\sum_{i\neq j}\dfrac{\hat{\sigma}^{z}_{i}\hat{\sigma}^{z}_{j}}{d(i,j)^{\alpha}}$ (71)
@@
Which is essentially the spin Hamiltonian (7) with an additional term (in the second line) describing the spin-boson coupling (with strength $\lambda$). The average dynamics of the system can be described by the Lindblad master equation,
$\dfrac{{\rm{d}}}{{\rm{d}}t}\hat{\rho}=-{\rm{i}}[\hat{H},\hat{\rho}]+\kappa\mathcal{D}[\hat{a}](\hat{\rho})\,,$ (72)
@@
${\rm{d}}\langle\hat{O}\rangle={}$ ${\rm{i}}{\rm{d}}t\mathopen{}\mathclose{{}\left\langle[\hat{H},\hat{O}]}\right\rangle+\kappa{\rm{d}}t\mathopen{}\mathclose{{}\left\langle\hat{a}^{\dagger}\hat{O}\hat{a}-\dfrac{1}{2}\mathopen{}\mathclose{{}\left\{\hat{a}^{\dagger}\hat{a},\hat{O}}\right\}}\right\rangle$ (73)
@@ D.1 Equations of motion
$\beta$ $\leftarrow 0\,,\quad$ (74)
@@
D.2 Results on the driven-dissipative long-range Tavis-Cummings model
Note that in the regime where the cavity dissipation is much faster than any other time scale in the system, i.e. when $\omega,J,\lambda\ll\kappa$, the cavity mode can be adiabatically eliminated and this model reduces effectively to the long-range spin model with collective spin dissipation $\hat{S}^{-}$ as considered in the main text, since the lossy cavity acts as a Markovian reservoir for the spin system that quickly evacuates entropy into the environment via the spin-boson interaction. To illustrate the generalization of our spin-wave quantum trajectory method to spin-boson systems, we place ourselves close to this regime, such that results qualitatively reminiscent of those in the spin-only model as presented in the main text can be expected. We fix the parameters as $s=1/2$, $\lambda=0.2\kappa$, $J=0.01\kappa$ and $\alpha=0.1$, and the results are shown in Fig. 10. The spin magnetization [panel (a)] displays similar behavior as compared to Fig. 5 (a), where a continuous dissipative phase transition predicted by the mean-field theory emerges. This transition is also accompanied by an entanglement phase transition [panel (b)] for the entanglement between the spins and the cavity, from an unentangled phase to an entangled one as the driving increases. The spin-wave density [panel (c)] displays a maximum close to the critical point, while the peak value decreases with the system size $N$, which is a signature of the long-range nature of the system. Finally, the spin-cavity entanglement also exhibits fast saturation, as shown in panel (d) for the case close to criticality. These results suggest a possible implementation of the long-range spin model with collective dissipation (as considered in the main text) in atom-cavity platforms, and that the brute-force experimental detection of the entanglement phase transition can also be realized with mitigated post-selection overhead.
Appendix E Comparison with other existing spin-wave methods
In this appendix, we provide a comparison between our method of spin-wave quantum trajectories (SWQT) and the method proposed in [97], which is a deterministic spin-wave method for dissipative systems (referred to as “DSW” in this section). The latter assumes the same approximations as in our method (i.e. lowest-order Holstein-Primakoff expansion and Gaussian approximation) except that their approximations are performed on the level of the averaged state (i.e. the density matrix) and their dynamical evolution is derived from the deterministic Lindblad master equation.
For illustration purposes, let us consider again the collective spin model defined in Sec. IV of the main text. The solutions given by the two methods are shown in Fig. 11 for comparison. Panel (a) shows the SWQT results for the parameters $\omega_{F}=1.25\kappa$, $J=0.1\kappa$, $S=64$, which is the same as that presented in Fig. 3 of the main text. The solution with the same parameters given by DSW is shown in panel (b), which is by no means close to the exact solution. As explained in the main text, this is a result of the highly mixed nature of the average state that the spin-wave approximations fail to capture. Panel (c) shows the spin-wave densities associated with the two solutions, where the SWQT has a spin-wave density that is smaller by several orders of magnitude, which is consistent with its accuracy compared to the exact solution.

### v2

@@ D.6 Benchmark of the binary approximation of the Gaussian noise
In this Section, we demonstrate a possible extension of our spin-wave theory, which allows the investigation of spin-boson systems. We adopt the same Gaussian approximation for the bosonic mode as for the bosonized spins, such that the state of the entire system can be characterized by the first and second moments, requiring a quadratic number of variables to specify the Gaussian variational ansatz in the general case. For illustration purposes, let us consider a driven-dissipative Tavis-Cummings model with long-range spin interactions. The spins are collectively driven and resonantly coupled to a single-mode cavity (with annihilation operator $\hat{a}$ for the cavity bosonic mode), and the cavity undergoes single-photon dissipation. The Hamiltonian can be written as follows,
@@
$\hat{H}=$ $\leavevmode\nobreak\ \omega\hat{S}^{x}+\dfrac{2sJ}{\mathcal{N}}\sum_{i\neq j}\dfrac{\hat{\sigma}^{z}_{i}\hat{\sigma}^{z}_{j}}{{\lVert\bm{r}_{i}-\bm{r}_{j}\rVert}^{\alpha}}$ (S16)
@@
which is essentially the spin Hamiltonian (7) in the main text with an additional term (in the second line) describing the spin-boson coupling (with strength $\lambda$). The average dynamics of the system can be described by the Lindblad master equation,
$\dfrac{{\rm{d}}}{{\rm{d}}t}\hat{\rho}=-{\rm{i}}[\hat{H},\hat{\rho}]+\kappa\mathcal{D}[\hat{a}](\hat{\rho})\,,$ (S17)
@@
${\rm{d}}\langle\hat{O}\rangle={}$ ${\rm{i}}{\rm{d}}t\mathopen{}\mathclose{{}\left\langle[\hat{H},\hat{O}]}\right\rangle+\kappa{\rm{d}}t\mathopen{}\mathclose{{}\left\langle\hat{a}^{\dagger}\hat{O}\hat{a}-\dfrac{1}{2}\mathopen{}\mathclose{{}\left\{\hat{a}^{\dagger}\hat{a},\hat{O}}\right\}}\right\rangle$ (S18)
@@ E.1 Equations of motion
$\beta$ $\leftarrow 0\,,\quad$ (S19)
@@
E.2 Results on the driven-dissipative long-range Tavis-Cummings model
Note that in the regime where the cavity dissipation is much faster than any other time scale in the system, i.e. when $\omega,J,\lambda\ll\kappa$, the cavity mode can be adiabatically eliminated and this model reduces effectively to the long-range spin model with collective spin dissipation $\hat{S}^{-}$ as considered in the main text, since the lossy cavity acts as a Markovian reservoir for the spin system that quickly evacuates entropy into the environment via the spin-boson interaction. To illustrate the generalization of our spin-wave quantum trajectory method to spin-boson systems, we place ourselves close to this regime, such that results qualitatively reminiscent of those in the spin-only model as presented in the main text can be expected. We fix the parameters as $s=1/2$, $\lambda=0.2\kappa$, $J=0.01\kappa$ and $\alpha=0.1$, and consider a 1D spin chain in the cavity. The results are shown in Fig. S10. The spin magnetization [panel (a)] displays similar behavior as compared to Fig. 3 (a) of the main text, where a continuous dissipative phase transition predicted by the mean-field theory emerges. This transition is also accompanied by an entanglement phase transition [panel (b)] for the entanglement between the spins and the cavity, from an unentangled phase to an entangled one as the driving increases. The spin-wave density [panel (c)] displays a maximum close to the critical point, while the peak value decreases with the system size $N$, which is a signature of the long-range nature of the system. Finally, the spin-cavity entanglement also exhibits fast saturation, as shown in panel (d) for the case close to criticality. These results suggest a possible implementation of the long-range spin model with collective dissipation (as considered in the main text) in atom-cavity platforms, and that the brute-force experimental detection of the entanglement phase transition can also be realized with mitigated post-selection overhead.